# Gemma: Open Models Based on Gemini Research and Technology

### v3

@@ 6.1 Human Preference Evaluations
On a held-out collection of around 1000 prompts oriented toward asking models to follow instructions across creative writing tasks, coding, and following instructions, Gemma 7B IT has a 51.7% positive win rate and Gemma 2B IT has a 41.6% win rate over Mistral v0.2 7B Instruct. On a held-out collection of around 400 prompts oriented towards testing basic safety protocols, Gemma 7B IT has a 58% win rate, while Gemma 2B IT has a 56.5% win rate. We report the corresponding numbers in Table 5.
@@ 6.2 Automated Benchmarks
SIQA 0-shot 48.3 50.3 $*{}^{*}$$*{}^{*}$ 49.7 51.8
Boolq 0-shot 77.4 81.7 $*{}^{*}$$*{}^{*}$ 69.4 83.2
@@
Table 6: Academic benchmark results, compared to similarly sized, openly-available models trained on general English text data. ${}^{\dagger}$ Mistral reports 50.2 on a different split for MBPP and on their split our 7B model achieves 54.5. $*{}^{*}$ evaluations run by us. Note that due to restrictive licensing, we were unable to run evals on LLaMA-2; all values above were previously reported in Touvron et al. (2023b).

### v4

@@ 6.1 Human Preference Evaluations
On a held-out collection of around 1000 prompts oriented toward asking models to follow instructions across creative writing tasks, coding, and following instructions, Gemma 7B IT has a 61.2% positive win rate and Gemma 2B IT has a 45% win rate over Mistral v0.2 7B Instruct. On a held-out collection of around 400 prompts oriented towards testing basic safety protocols, Gemma 7B IT has a 63.5% win rate, while Gemma 2B IT has a 60.1% win rate. We report the corresponding numbers in Table 5.
@@ 6.2 Automated Benchmarks
SIQA 0-shot 48.3 50.3 47.0 49.7 51.8
Boolq 0-shot 77.4 81.7 83.2 69.4 83.2
@@
Table 6: Academic benchmark results, compared to similarly sized, openly-available models trained on general English text data. Mistral reports 50.2 on a different split for MBPP and on their split our 7B model achieves 54.5. evaluations run by us. Note that due to restrictive licensing, we were unable to run evals on LLaMA-2; all values above were previously reported in Touvron et al. (2023b).